OpenStudy (anonymous):

Which of the following best describes the technique used to graph the equation using the slope and y-intercept? y = 5x + 10

OpenStudy (texaschic101):

can you tell from the equation y = 5x + 10, what the y intercept is ?

OpenStudy (texaschic101):

can you tell what the slope is ?

OpenStudy (texaschic101):

ok...in y = mx + b form, the number in the m position is the slope, and the number in the b position is the y intercept. y = 5x + 10 y = mx + b the number in the m position (the slope) is 5 the number in the b position (the y intercept) is 10....so it is (0,10) so graph your y intercept (0,10).....and since the slope is 5, or 5/1 if it makes it easier...go up 5 spaces (because 5 is positive) and to the right 1 space (because 1 is positive), then up 5, and to the right 1 and so on....